OpenStudy (anonymous):

Tristen and gorge are stuck simplifying radical expressions. Tristren has to simplify . William has to simplify . Using full sentences, describe how to fully simplify Vanessa and William's expressions. Describe if Tristena dn George started with equivalent expressions or if they started with expressions that are not equal?

OpenStudy (anonymous):

they are the same equations...

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (michele_laino):

Vanessa will simplify as below: \[{x^{4/3 - 5/6}} = {x^{3/6}} = {x^{1/2}} = \sqrt x \]

OpenStudy (michele_laino):

William will simplify as below: \[\sqrt[{16}]{{{x^{1 + 3 + 4}}}} = \sqrt[{16}]{{{x^8}}} = {x^{8/16}} = {x^{1/2}} = \sqrt x \]

OpenStudy (anonymous):

for vanessesa isn't it saying to divide?

OpenStudy (michele_laino):

yes! infact I have computed the difference between exponents

OpenStudy (anonymous):

okay:)

OpenStudy (anonymous):

Can you help with this one?

OpenStudy (michele_laino):

ok!

OpenStudy (michele_laino):

let's suppose that the graph of f(x) is: 

OpenStudy (michele_laino):

namely f(0)=5

OpenStudy (anonymous):

Yes

OpenStudy (michele_laino):

here is the graph of f(x)+2: